# Jet Tagging with More-Interaction Particle Transformer

### v2

@@ II.4 Implementation
The architecture of our MIParT model includes $K=5$ MI-particle attention blocks, $L=5$ particle attention blocks, and 2 class attention blocks. For particle embeddings $\mathbf{x}_{1}$, a three-layer Multi-Layer Perceptron (MLP) is used, with each layer containing 128, 512, and 64 neurons respectively. This configuration results in embeddings with a dimensionality of $D_{1}=64$. Each layer incorporates GELU as the activation function and Layer Normalization. Additionally, a three-layer, 64-channel Pointwise 1D Convolution is used for the interaction embeddings $\mathbf{U}_{1}$, performing convolutions only along the feature dimension. The $\mathbf{U}_{1}$ embeddings are further processed through a single-layer, 8-channel Pointwise 1D Convolution to generate $\mathbf{U}_{2}$, achieving a dimensionality of $D_{2}=8$. The MI-particle attention blocks implement MIA with 64 heads, while the P-MHA and Class Multi-Head Attention in the particle and class attention blocks utilize 8 heads each. A dropout rate of 0.1 is maintained in all MI-particle and particle attention blocks, with the class attention blocks being exempt from dropout.
@@ III Result and Discussion
In Table 5, we present comparative performance of various models on different sizes of the JetClass dataset. we displaced the results for the MIParT-L model alongside ParticleNet [34] and ParT [36] across 2M, 10M, and 100M JetClass datasets. We observe that as the dataset size increases, the performance of the models improves. Specifically, MIParT-L and ParT exhibit nearly identical effectiveness on very large datasets, surpassing that of ParticleNet.

### v3

@@ II.4 Implementation
The architecture of our MIParT model includes $K=5$ MI-particle attention blocks, $L=5$ particle attention blocks, and 2 class attention blocks. The choice of these hyperparameters balances complexity and accuracy; we observed an increase in accuracy with additional layers, but at the cost of increased complexity. Therefore, we limited the total number of attention blocks to ten. The rationale for choosing two class attention blocks follows the CaiT framework [38], which recommends such a configuration for efficient classification. For particle embeddings $\mathbf{x}_{1}$, a three-layer Multi-Layer Perceptron (MLP) is used, with each layer containing 128, 512, and 64 neurons respectively. This configuration results in embeddings with a dimensionality of $D_{1}=64$. The decision to reduce the embedding dimension compared to the ParT model was motivated by the addition of the MIA module. This adjustment allows us to rationalize the complexity of the model while maintaining its efficiency, thus optimizing the trade-off between performance and computational load. Each layer incorporates GELU as the activation function and Layer Normalization. Additionally, a three-layer, 64-channel Pointwise 1D Convolution is used for the interaction embeddings $\mathbf{U}_{1}$, performing convolutions only along the feature dimension. The $\mathbf{U}_{1}$ embeddings are further processed through a single-layer, 8-channel Pointwise 1D Convolution to generate $\mathbf{U}_{2}$, achieving a dimensionality of $D_{2}=8$. This design choice maintains consistency with the ParT model, ensuring alignment with established architectural standards and facilitating comparative analysis. The MI-particle attention blocks implement MIA with 64 heads, while the P-MHA and Class Multi-Head Attention in the particle and class attention blocks utilize 8 heads each. A dropout rate of 0.1 is maintained in all MI-particle and particle attention blocks, with the class attention blocks being exempt from dropout.
@@ III Result and Discussion
Given that MIParT shares many components with ParT and differs only in the addition of the MIA blocks, the comparative results between these two models highlight the effectiveness of the MIA block. Specifically, MIParT consists of 5 MIA blocks, 5 particle attention blocks, and 2 class attention blocks, whereas ParT consists of 8 particle attention blocks and 2 class attention blocks. Thus, from the results tested on the top tagging and quark-gluon datasets, it is evident that MIParT outperforms ParT, illustrating the significant role played by the MIA block. Furthermore, the effectiveness of the particle attention blocks has already been established in the ParT paper [36], and the impact of the class attention blocks has been tested in the CaiT framework [38].
Regarding the impact of hyperparameter choices on model performance, we find that MIParT is not overly sensitive to hyperparameter settings, but is more influenced by the overall network architecture. In particular, increasing the number of MIA blocks and particle attention blocks generally leads to better performance, but at the cost of increased complexity. Architectural modifications show that placing MIA blocks before particle attention blocks is optimal. Placing MIA blocks after particle attention blocks or alternating them significantly reduces effectiveness, sometimes to the point of performing worse than ParT. We think that MIA blocks function similarly to embeddings, allowing better integration of interaction information into the jets for improved information fusion and classification.
@@
In Table 5, we present comparative performance of various models on different sizes of the JetClass dataset. we displaced the results for the MIParT-L model alongside ParticleNet [34] and ParT [36] across 2M, 10M, and 100M JetClass datasets. We observe that as the dataset size increases, the performance of the models improves. Specifically, MIParT-L and ParT exhibit nearly identical effectiveness on very large datasets, surpassing that of ParticleNet. In addition, our evaluation of models on the JetClass dataset serves to test the ability of MIParT to generalize across different classification tasks. The JetClass datasets represent a more complex classification challenge, aimed at identifying Higgs boson decays to charm quarks. Our MIParT model shows remarkable stability on this task, highlighting its generalization capabilities.
Here, we discuss the improvements attributed to the pre-training performed on the JetClass dataset, with subsequent performance improvements observed on the top tagging and quark-gluon datasets. These three jet tagging tasks differ in their objectives: the JetClass dataset focuses on identifying Lorentz boosted $W$, $Z$, Higgs bosons and top quarks, the top tagging dataset aims to identify top quarks, and the quark-gluon dataset aims to distinguish between quark and gluon jets. The improvements across such diverse tasks suggest that MIParT has learned more generalized jet properties during the pre-training phase. These characteristics are effectively transferable to other tasks, demonstrating the model’s robustness and adaptability to different jet identification challenges. This capability highlights the potential of pre-trained models to improve performance in a wide range of applications by capturing and exploiting general features applicable to multiple scenarios.
Regarding the interpretability of MIParT, it is important to acknowledge that as a model based on the transformer neural network architecture, its interpretability remains limited, similar to many neural networks currently in use. Despite these interpretability challenges, the CMS collaboration has successfully used the graph neural network ParticleNet [34], another model that lacks full interpretability, to search for Higgs boson decay to charm quarks [49]. This success underscores that the lack of interpretability does not prevent the use of neural network models in particle physics experiments. In fact, ParticleNet, which functions as a non-interpretable “black box” model, has already begun to play a significant role in particle experiments, demonstrating that the non-interpretable nature of these models should not be a barrier to their use in advancing scientific discovery.